OpenStudy (anonymous):

A circle is defined by the following equation: -2x^(2)-2y^(2)+24x-54=0 What are the coordinates ofits center?

OpenStudy (anonymous):

(6,0)

OpenStudy (anonymous):

HEY CHIZZ whats d ans

OpenStudy (anonymous):

^what?

OpenStudy (anonymous):

first rewrite the equations as \[2x^2+2y^2-24x-54=0\] divide by 2 \[x^2+y^2-12x-27=0\] complete the square: \[(x-6)^2+y^2=27=36=63\] center is (6,0) radius is \[\sqrt{63}\]

OpenStudy (anonymous):

typo. 27+36 = 63

OpenStudy (anonymous):

another mistake. should be +54 not -54 so right hand side is -27+ 36 = 9 and radius is 3. sorry

OpenStudy (anonymous):

so is the center still (6,0)?